# Q-Adapter: Customizing Pre-trained LLMs to New Preferences with Forgetting Mitigation

### v3

@@ 5.1 Experimental Setup
Various methods for fine-tuning LLMs to specific downstream tasks have emerged recently, with some demonstrating potential in mitigating the the well-documented issue of forgetting. Specifically, we choose three representative categories of baselines to compare with Q-Adapter:
@@
We integrate the adapter module of Q-Adapter using LoRA (Hu et al., 2022), as our adapter is designed to complement the pre-trained model. This approach enhances training efficiency and significantly reduces inference costs, as the final policy is derived by combining the outputs of the base model, both with and without LoRA. To maintain consistency in trainable parameters, all baseline models also optimize parameters through a LoRA module. For the base model, we employ the Meta-Llama-3.1-8B-Instruct model, a state-of-the-art open-source LLM with superior benchmarking performance. Further details on the implementations of Q-Adapter and the baselines are provided in Appendix B.
Evaluation Metrics
  • Scores in popular benchmarks. Benchmarks including MMLU (Hendrycks et al., 2021b; a), MMLU Pro (Wang et al., 2024), GSM8k (Cobbe et al., 2021), BBH (Srivastava et al., 2022), and IFEval (Zhou et al., 2023) are selected, where MMLU and MMLU Pro measure the general knowledge of LLMs, GSM8k focuses on the math ability, BBH requires multi-reasoning, and IFEval evaluates the instruction following ability. Llama 3.1 models exhibit impressive performance on these benchmarks. Results on them are enough to reflect the general capabilities of the customized LLMs. We use lm-eval (Gao et al., 2023) to compute the results.
  • Judgments from LLMs. To evaluate how the customized LLM learns the new preference, we adopt a strong LLM like GPT-4o (OpenAI, 2023) as a judge. Specifically, we prompt GPT-4o to rank pairwise responses from SFT and the other methods, i.e., PR methods, Replay, and Q-Adapter. We thus use the win rate against SFT as our considered metric. We use alpaca-eval (Dubois et al., 2024) to compute the results.
It is worth noting that our results can be slightly different from those reported by the official Llama 3 paper due to different configurations. However, we make sure that comparisons are fair by evaluating all the methods under the same configuration. More details on evaluation are in Appendix C.
@@ 5.2 Customization to Domain-specific Datasets
Table 1: Evaluation on $r_{1}$ in DSP: general benchmark scores of Q-Adapter, SFT, Replay, PR (DPO), PR (PPO) fine-tuned in data from four specific domains in DSP. The scores represent accuracy in MMLU, MMLU Pro, flexible extraction accuracy in GSM8k, and average accuracy in IFEval over each level and matching.
@@
Table 2: Evaluation on $r_{2}$ in DSP: win rates of Q-Adapter, Replay, PR (DPO), PR (PPO) fine-tuned in data from four specific domains in DSP against the SFT baseline.
@@
Figure 2: Forgetting during training. Curves are the MMLU scores of Q-Adapter, Replay, and PR (DPO) under different training checkpoints.
We first consider customizing pre-trained LLMs in domain-specific datasets. The Domain-Specific Preference (DSP) dataset (Cheng et al., 2023) is a promising dataset that filters and categorizes data from the 52K Alpaca training set (Taori et al., 2023) into four domains including academy, business, entertainment, and literature. Each category of data contains different instructions and answers in its corresponding style, forming four distinct downstream tasks with different preferences. We post-train the Llama-3.1 model in data of each domain and evaluate the performance of all methods.
We use LLM benchmark scores to evaluate model capabilities in $r_{1}$ as a promising customized model should maintain its general performance without catastrophic forgetting. As shown in Table 1, we find that Q-Adapter generally outperforms other methods in these benchmark scores ranging from general knowledge, math, reasoning, and instruction following abilities. The results indicate that Q-Adapter effectively alleviates the forgetting issue. Moreover, replay-based methods and policy regularization methods also show moderate performance while SFT performs the worst due to the lack of an anti-forgetting mechanism. We also report the performance of the base model, where we find the performance drop of Q-Adapter in some benchmarks like MMLU, and BBH is less significant, well preserving the performance of the original model. To better illustrate the forgetting issue emerged in the fine-tuning process, we also evaluate the MMLU score of intermediate checkpoints in the academy domain data of DSP. As shown in Figure 2, we find that a significant performance drop can be observed in the training process of DPO and Replay. In contrast, Q-Adapter keeps a relatively stable performance in the MMLU metric due to its advantage of maximizing $r_{1}$ in the objective and preserving the pre-trained LLM $\pi_{1}^{*}$ for inference.
The second part of our evaluation is to assess the model capabilities in learning new preferences. As stated in Section 5.1, we adopt a strong LLM as the judgment to rank the generated responses of each method in the test dataset. Given an instruction, we first sample responses from all the methods. Then, we request LLM judgments through the OpenAI GPT-4o API to rank pairs of responses, one of which is from the model fine-tuned with SFT, while the other is from the model fine-tuned by Q-Adapter, PR or Replay. Finally, we calculate and compare the win rates of each method. To alleviate the preference of GPT-4o on long responses, we adopt a length-controlled win rates proposed in Dubois et al. (2024). The result is illustrated in Table 2. We can see that Q-Adapter achieves the best performance among all the four categories. This verifies the superior effectiveness of Q-Adapter in learning the new preference. We also showcase some generated responses in Appendix E.
@@ 5.3 Customization between Two Tasks
Figure 4: $r_{1}$ $r_{2}$ with different $\alpha_{0}$: Win rates of Q-Adapter against the SFT model with different choices of $\alpha_{0}$ in the helpful data ($r_{1}$) and the harmless data ($r_{2}$) from HH-RLHF.
We can further utilize the HH-RLHF dataset to investigate a key hyper-parameter $\alpha_{0}$, which balances between optimized rewards from the base model $r_{1}$ and those from the downstream task $r_{2}$ according to Equation 6. The hyper-parameter will also determine the inference policy. Therefore, we choose different values on $\alpha_{0}$ when customizing the model trained in helpful data to harmless data. Intuitively, a large $\alpha_{0}$ means more consideration on the base model, resulting in better performance in $r_{1}$, and vice versa. In Figure 4, we observe that a general tendency in obedience to this intuition occurs for $\alpha_{0}$ ranging from 0.03 to 0.001. An exception comes from a larger choice of $\alpha_{0}$ like 0.1, where we find training with large $\alpha_{0}$ can be unstable. We choose $\alpha_{0}=0.01$ for most of our experiments and note that the choice of $\alpha_{0}$ in a proper range will not be of great impact. We also conduct ablation studies to show the sensitivity of another hyper-parameter $\beta$ in Appendix D.
@@ 6 Related Work
Humans excel at understanding the intentions and emotions behind language. Based on grammatical syntax, common sense, social norms, etc., we can provide ranked preferences for different responses, which can serve as learning signals for language models and autonomous agents (Christiano et al., 2017). Ziegler et al. (2019) propose to first use the Bradley-Terry (BT) model (Bradley & Terry, 1952) to learn a reward function from human preferences and then apply modern RL algorithms, such as, PPO (Schulman et al., 2017), to fine-tune the LLM. This Reinforcement Learning from Human Feedback (RLHF) paradigm successfully aligns the LLM to human values and improves the model’s generalization (Kaufmann et al., 2023). Some of the most widely used chatbots, such as OpenAI’s ChatGPT (OpenAI, 2023) and Anthropic’s Claude (Anthropic, 2024), are trained by RLHF. To make RLHF more efficient, new variants have been proposed in subsequent works, including the replacement of PPO with REINFORCE (Williams, 1992; Li et al., 2024) and the derivation of new training objective by incorporating Kahnerman and Tversky’s prospect theory (Ethayarajh et al., 2024; Tversky & Kahneman, 1992). Our work differs from the above methods in two main ways: First, we do not learn the reward function; second, in order not to deviate too much from the pre-trained LLM, they use the KL-divergence constraint (Kullback & Leibler, 1951), while we restrict the customized LLM to maximize the reward used for aligning the pre-trained LLM. Although being effective, the RLHF pipeline is sophisticated, involving training multiple LLMs and sampling from the LLM in the loop of training. DPO (Rafailov et al., 2023), on the other hand, derives a simple classification loss but implicitly optimizes the same objective as existing RLHF algorithms, i.e., reward maximization with a KL-divergence constraint. Another method, IPL (Hejna & Sadigh, 2023), proposes an objective based on the Bellman operator (Puterman, 1994) and BT model to directly learn the Q-function from preference data. Despite holding a different motivation from ours and mainly focusing on control tasks, IPL inspires the technical derivation of our method.
@@
Currently, there is a growing number of open-source LLMs on the Internet (Wolf et al., 2019). However, most of these models are trained on generic datasets. To better apply pre-trained LLMs to downstream tasks, we need to further adapt them (Kirk et al., 2024). As fully fine-tuning the LLMs may consume a significantly huge amount of computation resources, Hu et al. (2022) propose LoRA which freezes the pre-trained model weights and adds trainable low-rank matrices into each layer of the Transformer architecture (Vaswani et al., 2017). In our practical implementation, we also use LoRA (please refer to Appendix B for more details). A commonly known issue with customization is that the LLM may forget its original knowledge and abilities. To alleviate this, a number of efforts have proposed different solutions. Ouyang et al. (2022) augment the reward with the KL-divergence between the customized and pre-trained LLMs. However, since the constraint is imposed only on the training data, the LLM can still forget a large part of its knowledge. Our approach is different in that we keep the pre-trained model intact and involve the pre-trained model in the inference process, ensuring that its knowledge is retained as much as possible. Huang et al. (2024) propose to generate data used to train the original LLM and replay it during customization. Nevertheless, the generated data may contain low-quality or unsafe responses, and training on them will exacerbate the LLM. Lin et al. (2024) try to find optimal combination ratios for each model layer and LoRA layer. Nevertheless, with the ratios being solved on only a small dataset, they are prone to converging to local optimums.
@@ 7 Conclusion
We propose Q-Adapter, which learns the residual Q-function directly from the data of the new human preference without learning a reward function. With the residual Q-function and the pre-trained LLM, we are able to generate responses that not only meet the preference in the dataset, but also inherit knowledge of the pre-trained LLM. The forgetting issue thus can be effectively alleviated, which was empirically verified by our experiments. In this age of increasingly open source LLMs, we believe that our work can better enable pre-trained LLMs to be applied to more specialized domains and tasks, while saving the cost of training LLMs from scratch.
@@
We acknowledge that this work has limitations. To derive the objective Equation 10 of Q-Adapter, we assume that the LLM to be customized has been pre-trained with RLHF. However, there are many open-source LLMs that are trained using other methods such as SFT. Investigating how these models can be effectively customized is also an interesting direction to pursue. Moreover, in our experiments, we only performed one round of customization. But our framework also supports continuously customizing the LLM, which can be achieved by learning more adapters. Since the main focus of this paper is to effectively customize a pre-trained LLM with forgetting mitigation, we plan to leave the directions mentioned above as future work.
@@ Acknowledgements
The authors would like to thank Peng-Yuan Wang, Ningjing Chao, Zhilong Zhang, Jing-Cheng Pang, Feng Chen, and Chen-Xiao Gao for their support and helpful discussions on improving the paper.
@@ Appendix A Proofs
$\max_{\pi}$ $))]\displaystyle\mathbb{E}_{s_{1}\sim\rho}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})-\alpha D_{\mathrm{KL}}\pqty{\pi(\cdot|s_{t})\|\pi_{\text{ref}}(\cdot|s_{t})}}}$
::=:= $])]\displaystyle\mathbb{E}_{s_{1}\sim\rho}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})-\alpha\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}\bqty{\log\frac{\pi(a_{t}|s_{t})}{\pi_{\text{ref}}(a_{t}|s_{t})}}}}$
$=$ $])]\displaystyle\mathbb{E}_{s_{1}\sim\rho}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})+\alpha\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}\bqty{\log\pi_{\text{ref}}(a_{t}|s_{t})}-\alpha\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}\bqty{\log\pi(a_{t}|s_{t})}}}$ ]
::=:= $)]\displaystyle\mathbb{E}_{s_{1}\sim\rho}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}^{\text{KL}}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))}}$
::=:= $\mathbb{E}_{s_{1}\sim\rho}\bqty{\mathbb{E}_{a_{1}\sim\pi(\cdot|s_{1})}Q^{\pi}(s_{1},a_{1})+\alpha\mathcal{H}(\pi(\cdot|s_{1}))},$
which concludes the proof. ∎
@@
Given state $s_{t}\in\mathcal{S}$ and action $a_{t}\in\mathcal{V}$, $\pi^{*}$ and its soft Q-function $Q^{\pi^{*}}$ satisfies
@@
$)].Q^{\pi^{*}}(s_{t},a_{t})=r_{\phi}^{\text{KL}}(s_{t},a_{t})+\gamma\mathbb{E}_{s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})}\bqty{\alpha\log\sum_{a_{t+1}\in\mathcal{V}}\exp\pqty{\frac{1}{\alpha}Q^{\pi^{*}}(s_{t+1},a_{t+1})}}.$ (12)
@@
$\tilde{\pi}^{*}(a|s)$ =$\overset{(a)}{=}\frac{\exp\pqty{\frac{1}{\tilde{\alpha}}\tilde{Q}(s,a)}}{\sum_{a^{\prime}\in\mathcal{V}}\exp\pqty{\frac{1}{\tilde{\alpha}}\tilde{Q}(s,a^{\prime})}}$
@@
where $(a)$ and $(c)$ are due to Equation 11, and $(b)$ comes from the definition of $\hat{Q}$. Moreover, by definition, we have that
$\hat{Q}(s,a)=$ $\tilde{Q}(s,a)-\lambda Q_{1}^{*}(s,a)$
@@
where $(a)$ and $(b)$ are due to Equation 12, and $(c)$ holds because of Equation 11. Notice that if we add $\lambda\alpha_{1}\log\pi_{1}(a|s)$ to both sides of Equation 5, we get
$Q^{\text{aug}}(s,a)\leftarrow r^{\text{aug}}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\bqty{\tilde{\alpha}\log\sum_{a^{\prime}\in\mathcal{V}}\exp\pqty{\frac{1}{\tilde{\alpha}}\pqty{Q^{\text{aug}}(s,a)}}},$
where $Q^{\text{aug}}(s,a)=Q(s,a)+\lambda\alpha_{1}\log\pi_{1}(a|s)$ and $r^{\text{aug}}(s,a)=r_{2}(s,a)+\lambda\alpha_{1}\log\pi_{1}(a|s)$. This is factually the soft Q-iteration on the reward function $r^{\text{aug}}$, whose convergence has already been proven in existing studies (see Theorem 3 of (Haarnoja et al., 2017)). ∎
Appendix B Implementations Details
B.1 Base Model
We use a base model of Meta-Llama-3.1-8B-Instruct (Dubey et al., 2024) from Meta in the DSP tasks. The model is directly retrieved from Hugging Face [^5] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. The 8B version model well balances between effectiveness and efficiency and is convenient to be fine-tuned with LoRA (Hu et al., 2022). After pre-trained in a plethora of data and, the instruct version of Llama-3.1 8B models is further post-trained with SFT and DPO to enhance its instruction following ability, which also makes it satisfy the conditions of a base model in Q-Adapter, i.e., well aligned with some values via RL. When adding LoRA layers over the base model, we adopt the peft library (Mangrulkar et al., 2022) to modify the linear layers of the model with a rank of $8$. Other LoRA parameters, including a scaling of $16$ and a dropout probability of $0.05$, follow the default configuration. When performing fine-tuning with LoRA, we quantitize the parameters of the base model to 8-bit integer format. The number of trainable parameters during the post-training process is about 3M.

### v4

@@ 5.1 Experimental Setup
Datasets
To compare the performance of Q-Adapter with baselines in customizing pre-trained LLMs, we consider two scenarios: one where we would like to customize a general-purpose LLM with domain-specific datasets and the other where we consider customizing a model learned in one domain to another. For the first one, we choose the Domain Specific Preference (DSP) (Cheng et al., 2023) dataset that contains preference data from four domains including academy, business, entertainment and literature. For the second one, we choose the HH-RLHF (Bai et al., 2022) dataset that has two folds of preference data: helpful and harmless. The helpful data will first be used to post-train a Llama-3.1 model. Then we will use the harmless data to further customize it.
Evaluation Metrics
  • Scores in popular benchmarks. Benchmarks including MMLU (Hendrycks et al., 2021b; a), MMLU Pro (Wang et al., 2024), GSM8k (Cobbe et al., 2021), BBH (Srivastava et al., 2022), and IFEval (Zhou et al., 2023) are selected to reflect the general capabilities of the customized LLMs. Here, MMLU and MMLU Pro measure the general knowledge of LLMs, GSM8k focuses on the math ability, BBH requires multi-reasoning, and IFEval evaluates the instruction following ability.
  • Judgments from LLMs. To evaluate how the customized LLM learns the new preference, we adopt a strong LLM like GPT-4o (OpenAI, 2023) as a judge. Specifically, we prompt GPT-4o to rank pairwise responses from SFT and the other methods, i.e., PR methods, Replay, and Q-Adapter. We thus use the win rate against SFT as our considered metric. We use AlpacaEval (Dubois et al., 2024) for the auto-evaluation process.
@@
Various methods for fine-tuning LLMs to specific downstream tasks have emerged recently, with some demonstrating potential in mitigating the well-documented issue of forgetting. Specifically, we choose three representative categories of baselines to compare with Q-Adapter:
@@
We integrate the adapter module of Q-Adapter using LoRA (Hu et al., 2022), as our adapter is designed to complement the pre-trained model. This approach enhances training efficiency and significantly reduces inference costs, as the final policy is derived by combining the outputs of the base model, both with and without LoRA. To maintain consistency in trainable parameters, all baseline models also optimize parameters through a LoRA module. For the base model, we employ the Meta-Llama-3.1-8B-Instruct model (Team, 2024b), a state-of-the-art open-source LLM with superior benchmarking performance. Further details on the implementations of Q-Adapter and the baselines are provided in Appendix C. Note that our results of the base model can be slightly different from those reported by the official Llama 3 paper (Team, 2024b) due to different configurations. However, we ensure that comparisons are fair by evaluating all the methods under the same configuration. More details on evaluation are in Appendix D.
@@ 5.2 Customization to Domain-specific Datasets
We first consider customizing pre-trained LLMs in domain-specific datasets. The Domain-Specific Preference (DSP) dataset (Cheng et al., 2023) is a promising dataset that filters and categorizes data from the 52K Alpaca training set (Taori et al., 2023) into four domains including academy, business, entertainment, and literature. Each category of data contains different instructions and answers in its corresponding style, forming four distinct downstream tasks with different preferences. We post-train the Llama-3.1 model in data from each domain and evaluate the performance of all methods.
Table 1: Evaluation on $r_{1}$ in DSP: general benchmark scores of Q-Adapter, SFT, Replay, PR (DPO), and PR (PPO) fine-tuned in data from four specific domains in DSP. The scores represent accuracy in MMLU, MMLU Pro, flexible extraction accuracy in GSM8k, and average accuracy in IFEval over each level and matching.
@@
Table 2: Evaluation on $r_{2}$ in DSP: win rates of Q-Adapter, Replay, PR (DPO), and PR (PPO) fine-tuned in data from four specific domains in DSP against the SFT baseline.
@@
We use LLM benchmark scores to evaluate model capabilities in $r_{1}$ as a promising customized model should maintain its general performance without catastrophic forgetting. As shown in Table 1, we find that Q-Adapter generally outperforms other methods in these benchmark scores ranging from general knowledge, math, reasoning, and instruction following abilities. The results indicate that Q-Adapter effectively alleviates the forgetting issue. Moreover, replay-based methods and policy regularization methods also show moderate performance while SFT performs the worst due to the lack of an anti-forgetting mechanism. We also report the performance of the base model, where we find the performance drop of Q-Adapter in some benchmarks like MMLU, and BBH is less significant, well preserving the performance of the original model. The Base Model performs best, as there is some forgetting during the customization process. To better illustrate the forgetting issue emerging in the fine-tuning process, we evaluate the MMLU score of different intermediate checkpoints in the academy domain data of DSP. As shown in Figure 3, we find that a significant performance drop can be observed in the training processes of DPO and Replay. In contrast, Q-Adapter keeps a relatively stable performance in the MMLU metric due to its advantage of maximizing $r_{1}$ in the objective and preserving the pre-trained LLM $\pi_{1}^{*}$ for inference.
The second part of our evaluation is to assess the model capabilities in learning new preferences. As stated in Section 5.1, we adopt a strong LLM as the judgment to rank the generated responses of each method in the test dataset. Given an instruction, we first sample responses from all the methods. Then, we request LLM judgments through the OpenAI GPT-4o API to rank pairs of responses, one of which is from the model fine-tuned with SFT, while the other is from the model fine-tuned by Q-Adapter, PR or Replay. Finally, we calculate and compare the win rates of each method. To alleviate the preference of GPT-4o on long responses, we adopt a length-controlled win rates proposed in Dubois et al. (2024). The result is illustrated in Table 2. We can see that Q-Adapter achieves the best performance among all the four categories. This verifies the superior effectiveness of Q-Adapter in learning the new preference. We also showcase some generated responses in Appendix E. Meanwhile, potential model bias can be introduced when using GPT-4o as the evaluator. To ablate this, we also re-evaluate the generation texts with Claude-3.5-Sonnet (Anthropic, 2024) and Deepseek-chat (Liu et al., 2024a) and find consistent results. The details can be found in Section D.2.
@@ 5.3 Customization between Two Tasks
Figure 2: Visualization of forgetting during training. The curves illustrate the MMLU scores of Q-Adapter, Replay, and PR (DPO) under different training checkpoints.
@@ 6 Related Work
Humans excel at understanding the intentions and emotions behind language. Based on grammatical syntax, common sense, social norms, etc., we can provide ranked preferences for different responses, which can serve as learning signals for language models and autonomous agents (Christiano et al., 2017). Ziegler et al. (2019) propose to first use the Bradley-Terry (BT) model (Bradley & Terry, 1952) to learn a reward function from human preferences and then apply modern RL algorithms, such as PPO (Schulman et al., 2017), to fine-tune the LLM. This Reinforcement Learning from Human Feedback (RLHF) paradigm successfully aligns the LLM to human values and improves the model’s generalization (Kaufmann et al., 2023). Some of the most widely used chatbots, such as OpenAI’s ChatGPT (OpenAI, 2023) and Anthropic’s Claude (Anthropic, 2024), are trained by RLHF. To make RLHF more efficient, new variants have been proposed in subsequent works, including the replacement of PPO with REINFORCE (Williams, 1992; Li et al., 2024) and the derivation of new training objective by incorporating Kahnerman and Tversky’s prospect theory (Ethayarajh et al., 2024; Tversky & Kahneman, 1992). Our work differs from the above methods in two main ways: First, we do not learn the reward function; second, in order not to deviate too much from the pre-trained LLM, they use the KL-divergence constraint (Kullback & Leibler, 1951), while we restrict the customized LLM to maximize the expected discounted sum of the reward used for aligning the pre-trained LLM. Although being effective, the RLHF pipeline is sophisticated, involving training multiple LLMs and sampling from the LLM in the loop of training. DPO (Rafailov et al., 2023), on the other hand, derives a simple classification loss but implicitly optimizes the same objective as existing RLHF algorithms, i.e., reward maximization with a KL-divergence constraint. Another method, IPL (Hejna & Sadigh, 2023), proposes an objective based on the Bellman operator (Puterman, 1994) and BT model to directly learn the Q-function from preference data. Despite holding a different motivation from ours and mainly focusing on control tasks, IPL inspires the technical derivation of our method.
@@
Currently, there is a growing number of open-source LLMs on the Internet (Zhao et al., 2024). However, most of these models are trained on generic datasets. To better apply pre-trained LLMs to downstream tasks, we need to further adapt them (Kirk et al., 2024). As fully fine-tuning the LLMs may consume a significantly huge amount of computational resources, Hu et al. (2022) propose LoRA which freezes the pre-trained model weights and adds trainable low-rank matrices into each layer of the Transformer architecture (Vaswani et al., 2017). In our practical implementation, we also use LoRA (please refer to Appendix C for more details). A commonly known issue with customization is that the LLM may forget its original knowledge and abilities. To alleviate this, a number of efforts have proposed different solutions. Ouyang et al. (2022) augment the reward with the KL-divergence between the customized and pre-trained LLMs. However, since the constraint is imposed only on the training data, the LLM can still forget a large part of its knowledge. Our approach is different in that we keep the pre-trained model intact and involve the pre-trained model in the inference process, ensuring that its knowledge is retained as much as possible. Huang et al. (2024) propose to generate data used to train the original LLM and replay it during customization. Nevertheless, the generated data may contain low-quality or unsafe responses, and training on them will exacerbate the LLM. Lin et al. (2024) try to find optimal combination ratios for each model layer and LoRA layer. Nevertheless, with the ratios being solved on only a small dataset, they are prone to converging to local optima.
@@ 7 Conclusion
We propose Q-Adapter, which learns the residual Q-function directly from the data of the new human preference without learning a reward function. With the residual Q-function and the pre-trained LLM, we are able to generate responses that not only meet the preference in the dataset, but also inherit knowledge of the pre-trained LLM. The forgetting issue thus can be effectively alleviated, which was empirically verified by our experiments. In this age of increasingly open-source LLMs, we believe that our work can better enable pre-trained LLMs to be applied to more specialized domains and tasks (Chen et al., 2024), while saving the cost of training from scratch.
@@
Our work has several limitations. To derive the objective of Q-Adapter in Equation 12, we assume that the LLM to be customized has been pre-trained with RLHF. However, there are many open-source LLMs that are trained using other methods such as SFT. Investigating how these models can be effectively customized is also an interesting direction to pursue. Moreover, in our experiments, we only performed one round of customization. But our framework also supports continuously customizing the LLM, which can be achieved by learning more adapters. Additionally, although the derivation of Q-Adapter does not require high quality and particular relevance of the training preference data, bias and noise from data may affect the model performance. Further investigating how these factors affect LLM customization is an interesting topic. Last but not least, we have observed that in practical tasks, the collaboration of multiple LLMs may be involved. Using multi-agent reinforcement learning methods (Yuan et al., 2023a) to customize a group of LLMs is also an interesting topic. We plan to leave the directions mentioned above as future work.
@@ Acknowledgements
This work is supported by the National Science Foundation of China (62276126, 62250069, U24A20324) and the Natural Science Foundation of Jiangsu (BK2024119). The authors would like to thank Peng-Yuan Wang, Xuqin Zhang, Ningjing Chao, and Jing-Cheng Pang for their support and helpful discussions on improving the paper.
@@ Appendix A Proofs
$\max_{\pi}$ $))]\displaystyle\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi(\cdot|s_{t})}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})-\alpha D_{\mathrm{KL}}\pqty{\pi(\cdot|s_{t})\|\pi_{\mathrm{ref}}(\cdot|s_{t})}}}$
$=$ $])]\displaystyle\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi(\cdot|s_{t})}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})-\alpha\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}\bqty{\log\frac{\pi(a_{t}|s_{t})}{\pi_{\mathrm{ref}}(a_{t}|s_{t})}}}}$
$=$ $])]\displaystyle\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi(\cdot|s_{t})}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}(s_{t},a_{t})+\alpha\log\pi_{\mathrm{ref}}(a_{t}|s_{t})-\alpha\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}\bqty{\log\pi(a_{t}|s_{t})}}}$
$=$ $)]\displaystyle\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi(\cdot|s_{t})}\bqty{\sum_{t=1}^{T}\gamma^{t-1}\pqty{r_{\phi}^{\mathrm{KL}}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))}}$
$=$ $\mathbb{E}_{s_{1}\sim\rho}\bqty{\mathbb{E}_{a_{1}\sim\pi(\cdot|s_{1})}Q^{\pi}(s_{1},a_{1})+\alpha\mathcal{H}(\pi(\cdot|s_{1}))},$
where the last step follows from the definition of $Q^{\pi}$, as in Equation 4. ∎
@@
Given state $s_{t}\in\mathcal{S}$ and action $a_{t}\in\mathcal{V}$, $\pi^{*}$ and its soft Q-function $Q^{\pi^{*}}$ satisfy
@@
$)].Q^{\pi^{*}}(s_{t},a_{t})=r_{\phi}^{\mathrm{KL}}(s_{t},a_{t})+\gamma\mathbb{E}_{s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})}\bqty{\alpha\log\sum_{a_{t+1}\in\mathcal{V}}\exp\pqty{\frac{1}{\alpha}Q^{\pi^{*}}(s_{t+1},a_{t+1})}}.$ (14)
@@
$\tilde{\pi}^{*}(a|s)$ =$\overset{(a)}{=}\frac{\exp\pqty{\frac{1}{\tilde{\alpha}}\tilde{Q}^{*}(s,a)}}{\sum_{a^{\prime}\in\mathcal{V}}\exp\pqty{\frac{1}{\tilde{\alpha}}\tilde{Q}^{*}(s,a^{\prime})}}$
@@
where $(a)$ and $(c)$ are due to Equation 13, and $(b)$ comes from Equation 5. Moreover, by definition, we have that
$\hat{Q}(s,a)=$ $\tilde{Q}^{*}(s,a)-\lambda Q_{1}^{*}(s,a)$
@@
where $(a)$ and $(b)$ are due to Equation 14, and $(c)$ holds because of Equation 5 and Equation 13. Notice that if we add $\lambda\alpha_{1}\log\pi^{*}_{1}(a|s)$ to both sides of Equation 7, we get
$Q^{\text{aug}}(s,a)\leftarrow r^{\text{aug}}(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\bqty{\tilde{\alpha}\log\sum_{a^{\prime}\in\mathcal{V}}\exp\pqty{\frac{1}{\tilde{\alpha}}\pqty{Q^{\text{aug}}(s^{\prime},a^{\prime})}}},$
where $Q^{\text{aug}}(s,a)=Q(s,a)+\lambda\alpha_{1}\log\pi^{*}_{1}(a|s)$ and $r^{\text{aug}}(s,a)=r_{2}(s,a)+\lambda\alpha_{1}\log\pi^{*}_{1}(a|s)$. This is factually the soft Q-iteration on the reward function $r^{\text{aug}}$, whose convergence has already been proven in existing studies (see Theorem 3 of (Haarnoja et al., 2017)). ∎
Appendix B Sensitivity Analysis
We further utilize the HH-RLHF dataset to investigate a key hyper-parameter $\alpha_{0}$, which balances between optimized rewards from the base model $r_{1}$ and those from the downstream task $r_{2}$ according to Equation 8. The hyper-parameter will also determine the inference policy. Therefore, we choose different values on $\alpha_{0}$ when customizing the model trained in helpful data to harmless data. Intuitively, a large $\alpha_{0}$ means more consideration on the base model, resulting in better performance in $r_{1}$, and vice versa. In Figure 5, we observe that a general tendency in obedience to this intuition occurs for $\alpha_{0}$ ranging from 0.03 to 0.001. An exception comes from a larger choice of $\alpha_{0}$ like 0.1, where we find training with large $\alpha_{0}$ can be unstable. We choose $\alpha_{0}=0.01$ for most of our experiments and note that the choice of $\alpha_{0}$ in a proper range will not be of great impact.
In Equation 12, we add an additional regularization loss term in an L2 form to constrain the estimated reward range with a hyper-parameter $\beta$. We find that adding such a regularization term can stabilize the training process and thus result in better performance. In Figure 5, we show the MMLU scores of Q-Adapter with different choices of $\beta$. The results demonstrate that introducing the regularization leads to better performance in MMLU. On the other hand, the value of $\beta$ is not very sensitive from $0.01$ to $1.0$. In our main experiments, we train Q-Adapter with $\beta=0.1$.
Figure 4: Win rates of Q-Adapter against the SFT model with different $\alpha_{0}$ on HH-RLHF.
Appendix C Implementations Details
C.1 Base Model
We use a base model of Meta-Llama-3.1-8B-Instruct (Team, 2024b) from Meta in the DSP tasks. The model is directly retrieved from Hugging Face [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. The 8B version model well balances between effectiveness and efficiency and is convenient to be fine-tuned with LoRA (Hu et al., 2022). After pre-trained in a plethora of data and, the instruct version of Llama-3.1 8B models is further post-trained with SFT and DPO to enhance its instruction following ability, which also makes it satisfy the conditions of a base model in Q-Adapter, i.e., well aligned with some values via RL. When adding LoRA layers over the base model, we adopt the peft library (Mangrulkar et al., 2022) to modify the linear layers of the model with a rank of $8$. Other LoRA parameters, including a scaling of $16$ and a dropout probability of $0.05$, follow the default configuration. When performing fine-tuning with LoRA, we quantitize the parameters of the base model to 8-bit integer format. The number of trainable parameters during the post-training process is about 3M.